# Blending Data-Driven Priors in Dynamic Games

### v2

@@ IV-C1 The LQL approximation
Backward Pass. During the backward pass, we linearize dynamics (1) and quadraticize cost (2) around ${\eta}$ to obtain a linear dynamical system and quadratic cost functions that fit into the assumption of KL-LQG game (c.f. Theorem 1). Next, we use the Laplace approximation [85, Chapter 4] to obtain a Gaussian distribution that captures the reference policy locally around ${\eta}$. Specifically, the conditional probability distribution of a reference policy ${\tilde{\pi}}^{i}_{t}({u}^{i}_{t}|{x}_{t})$ is approximated as:
@@
Line Search. A common issue in ILQ-based trajectory optimization is that directly applying the mean of the KL-LQG strategy in the forward pass may cause the system to deviate too much from the nominal trajectory ${\eta}$, resulting in divergence of the algorithm. A possible remedy is line search [86, Chapter 3], which controls the trajectory update by scaling down the KL-LQG strategy parameters. Specifically, for a line search parameter $\varepsilon\in(0,1]$ and KL-LQG strategy $\mathcal{N}(-K^{i}_{t}{x}_{t}-\kappa^{i}_{t},\Sigma_{t}^{i})$, the adjusted strategy is $\mathcal{N}(-\varepsilon K^{i}_{t}{x}_{t}-\varepsilon\kappa^{i}_{t},\varepsilon\Sigma_{t}^{i})$, which is then used in the forward pass to obtain a new nominal trajectory ${\eta}$. The procedure starts with $\varepsilon\leftarrow 1$ and repeats by decreasing $\varepsilon$ by half until a convergence criterion is met. Two commonly used criteria are decreasing of the social cost (i.e., the sum of all players’ costs) and that the updated nominal trajectory is sufficiently close to the previous one.
@@ V-A Two-Player Tollbooth Selection
Figure 3: Two-player tollbooth interaction. Player 1 (red) and Player 2 (blue) are approaching a tollbooth on a twoilane road (upper: 1, lower: 2). Both players receive a coordination bonus for taking opposite lanes (e.g. for reducing traffic), but Player 2 has a strong preference to merge into Lane 1. Since players incur a large penalty for deviating from the lane centerline, Player 1 may get caught in a local cost minimum (Lane 1) without external guidance to switch to Lane 2.
@@
ILQGames $0.00$ $\mathbf{1.00}$ $45.0$ $9.50$ $8.45$
MaxEnt $0.28$ $0.68$ $46.8$ $6.43$ $6.91$
KLGame (Ours) $\mathbf{1.00}$ $\mathbf{1.00}$ $\mathbf{51.8}$ $8.18$ 4.22
@@
Table I presents performance comparisons for the three methods over $100$ trials. We report the following methods: (i) coordination rate (CR): the proportion of trials that end with the vehicles in opposite lanes, (ii) safety rate (SR): the proportion of trials where the vehicles remain in-bounds and do not collide, (iii) road progress, the distance that Player 1 travels from their initial position (as a measure of efficiency), (iv) the minimum distance to Player 2, and (v) the time-averaged cost incurred over the full rollout, averaged over all of the trials. Using the a simple reference policy, KLGame is able to coordinate at a much higher rate than other methods while maintaining safety.
@@ V-B Autonomous Car Racing
Reference policy and rival behaviors. To obtain a reference policy, we use a driving simulator (Fig. 6) to collect driving data (states, actions, track information, etc.) from races performed by two expert human racers in Carla [90]. With this data, we employ an inverse game training procedure inspired by works such as [37, 38, 17]. This approach involves defining a set of basis functions for the game’s cost—each representing a distinct racing specification like optimizing lap times, collision avoidance, and track boundary adherence—and learning a context-dependent cost weight model. The model, which takes as inputs driving data and outputs the cost weights, is learned as a deep neural network leveraging automatic differentiation in JAX [89]. Given the limited volume of human-generated data, we use a dataset aggregation strategy [91] to enrich the training dataset for the reference policy. To obtain a stochastic reference policy, we use a game-theoretic variant of the Boltzmann distribution [9, Section 3.2], in which the Q-value function is obtained by solving the game with the inverse-learned cost weight model, which captures the noisily-rational aspect of human decision making. Finally, we produce the rival’s behavior with a defensive ILQGame policy [5], whose parameters (cost terms and weights) are not accessible to either the reference or KLGame policies.
@@
Multi-modal KLGame. Motivated by the result shown above for Basic KLGame, where blending only an aggressive reference policy can lead to unsafe outcomes, this experiment demonstrates how a multi-modal KLGame (MM-KLGame) (c.f. Sec. IV-C2) with an additional safety-enhancing mode performs in racing scenario. The second mode we use is a following policy that controls the ego car to (i) follow the rival while keeping a constant distance when the rival is leading or (ii) defend against the rival after overtaking it. During each planning cycle, instead of picking the mode via sampling, we use rollout-based safety filters [92, Section 3.3] and perform a collision check for the KLGame-optimized trajectories associated with the overtaking mode: if they are collision-free, we choose the overtaking mode; otherwise, we pick the following mode. Again, we vary the value of $\lambda$ and inspect the qualitative difference in the resulting ego behaviors. The simulation snapshots are shown in Fig. 8. When $\lambda$ takes a relatively small value, $\lambda=5$, the MM-KLGame is not sufficiently guided by the overtaking policy and fails to overtake the rival. When $\lambda=50$, the ego using MM-KLGame behaves similarly to the unimodal case and safely overtakes the rival. However, in the case when $\lambda=5\times 10^{4}$, instead of incurring a collision as in the unimodal KLGame case, the MM-KLGame leads to a safe overtake thanks to the additional following mode in the multi-modal reference policy. We also apply the reference policy to the ego using the same mode selection rule. Interestingly, the ego car fails to overtake the rival due to a more frequent application of the following policy triggered by the unsafe overtaking policy, which ultimately slows down the ego vehicle. This example also demonstrates the potential of KLGame using a multi-modal reference policy that mixes a data-driven policy and an optimization-based policy.
@@ V-C Scene-Consistent Simulated Agents in Waymax
KLGame on Waymax. For our implementation of KLGame on Waymax, we use only two generic optimization targets as cost functions: collision distance and control magnitude regularization, with the rest of the agent’s optimization objective coming from the MTR reference policy guidance. Alongside the control regularization, each agent is also tasked with maximizing its signed distance to the closest players and off-road areas. These objectives prompt the agents to prevent collisions effectively and adhere to the road. The signed distance between agents is calculated as the Minkowski difference between two footprints, while the distance to the off-road areas is obtained from road boundaries. In our experiments, all agents are replanned every second in a receding horizon fashion with a 2-seconds planning horizon and 0.1s discrete timestep.
Figure 9: As car 1 pulls out from the parked vehicle, KLGame optimizes the joint trajectories, guiding both agents 0 and 1 to exhibit yielding behaviors to prevent a collision. However, the reference MTR policy, in contrast, results in a collision.
@@
Finally, Fig. 11 shows that KLGame can yield diverse, yet scene-consistent traffic plans, particularly when paired with a multi-modal reference policy. In this intersection scenario, agent 44 approaches the intersection from the wrong side of the road. This unlikely event leads to the marginal predictions from MTR having a hard time keeping the scenario collision-free. In the first rollout, as vehicle 44 opts for a mode to proceed straight, KLGame plans emergency maneuvers for both vehicles to circumvent collision. In the subsequent rollout, vehicle 44 samples the mode of turning right. Despite a proximate interaction, KLGame computes the optimal joint trajectory, thereby successfully preventing a collision.
Figure 10: As the agent 1 merges into the main road, the MTR-derived policy slows down vehicle 0 but ultimately results in a collision. In comparison, KLGame optimizes the joint trajectories by planning for vehicle 0 to switch lanes and yield, thus avoiding a collision.
Figure 11: Agent 44 approaches the intersection from the incorrect side of the road. In the first rollout, vehicle 44 opts for a mode to proceed straight, causing KLGame to plan an emergency maneuver to circumvent collision. In the subsequent rollout, vehicle 44 samples the mode of turning right. Despite a proximate interaction, KLGame computes the optimal joint trajectory, thereby successfully averting a collision. This demonstrates KLGame’s ability to generate diverse trajectories with a multi-modal reference policy.

### v3

@@ IV-C1 The LQL approximation
Backward Pass. During the backward pass, we linearize dynamics (1) and quadraticize cost (2) around ${\eta}$ to obtain a linear dynamical system and quadratic cost functions that fit into the assumption of KL-LQG game (c.f. Theorem 1). Next, we use the Laplace approximation [89, Chapter 4] to obtain a Gaussian distribution that captures the reference policy locally around ${\eta}$. Specifically, the conditional probability distribution of a reference policy ${\tilde{\pi}}^{i}_{t}({u}^{i}_{t}|{x}_{t})$ is approximated as:
@@
Line Search. A common issue in ILQ-based trajectory optimization is that directly applying the mean of the KL-LQG strategy in the forward pass may cause the system to deviate too much from the nominal trajectory ${\eta}$, resulting in divergence of the algorithm. A possible remedy is line search [90, Chapter 3], which controls the trajectory update by scaling down the KL-LQG strategy parameters. Specifically, for a line search parameter $\varepsilon\in(0,1]$ and KL-LQG strategy $\mathcal{N}(-K^{i}_{t}{x}_{t}-\kappa^{i}_{t},\Sigma_{t}^{i})$, the adjusted strategy is $\mathcal{N}(-\varepsilon K^{i}_{t}{x}_{t}-\varepsilon\kappa^{i}_{t},\varepsilon\Sigma_{t}^{i})$, which is then used in the forward pass to obtain a new nominal trajectory ${\eta}$. The procedure starts with $\varepsilon\leftarrow 1$ and repeats by decreasing $\varepsilon$ by half until a convergence criterion is met. Two commonly used criteria are decreasing of the social cost (i.e., the sum of all players’ costs) and that the updated nominal trajectory is sufficiently close to the previous one.
@@ IV-C2 Accounting for multi-modality with scenario optimization
Computation Complexity. The computation complexity of multi-modal KLGame is affected by the state dimension $n$, each player’s control dimension $m_{i}$, the number of players $N$, the number of scenario tree branches $H$, and the number of LQL iterations (c.f. Section IV-C1). Assuming $n>m_{i}$ for each player $i$, in each iteration, along a single tree branch, the complexity of the LQ approximation and solving the coupled Riccati equation of the corresponding KL-LQG game (c.f. [1, Corollary 6.1]) are $\mathcal{O}(Nn^{2})$ and $\mathcal{O}(N^{3}n^{3})$, respectively, coinciding with that of ILQGame [5]. The Laplace approximation has complexity $\mathcal{O}(Nn^{3})$. The computation bottleneck of Algorithm 1 is the matrix inversion operation when solving the coupled Riccati equation, which has complexity $\mathcal{O}(N^{3}n^{3})$. Therefore, the overall per-iteration complexity of multi-modal KLGame is $\mathcal{O}(HN^{3}n^{3})$ (for ILQGame, the complexity is $\mathcal{O}(N^{3}n^{3})$). Our implementation leverages the automatic vectorization of JAX [92] to efficiently batch-compute gradients and Hessians across different players and time steps. Throughout our experiments, we observe that Algorithm 1, when warmstarted and solved in a receding horizon fashion, typically converges with a small number of LQL iterations. As shown in Fig. 12, our solver achieves a 10Hz planning rate for a reasonable number of agents (around four) in complex traffic interaction scenarios where agents are in close proximity.
@@ V-A Two-Player Tollbooth Selection
ILQGames
MaxEnt
KLGame
Figure 3: Two-player tollbooth interaction. Player 1 (red) and Player 2 (blue) are approaching a tollbooth on a two-lane road (upper: 1, lower: 2). Both players receive a coordination bonus for taking opposite lanes (e.g. for reducing traffic), but Player 2 has a strong preference to merge into Lane 1. Since players incur a large penalty for deviating from the lane centerline, Player 1 may get caught in a local cost minimum (Lane 1) without external guidance to switch to Lane 2.
@@
ILQGames $00.00\pm 0$ 0.00 $\mathbf{1.00}\pm 0.00$ $45.0\pm 0.00$ $8.45\pm 0.00$
MaxEnt $0.28\pm 0.43$ $0.68\pm 0.48$ $46.8\pm 3.03$ $6.91\pm 1.64$
KLGame (Ours) $\mathbf{1.00}\pm 0.00$ $\mathbf{1.00}\pm 0.00$ $\mathbf{51.8}\pm 5.83$ $\textbf{4.22}\pm 1.12$
@@
Table I presents performance comparisons for the three methods over $100$ trials. We report the following methods: (i) coordination rate (CR): the proportion of trials that end with the vehicles in opposite lanes, (ii) safety rate (SR): the proportion of trials where the vehicles remain in-bounds and do not collide, (iii) road progress, the distance that Player 1 travels from their initial position (as a measure of efficiency), (iv) the minimum distance to Player 2, and (v) the time-averaged cost incurred over the full rollout, averaged over all of the trials. Using a simple reference policy, KLGame is able to coordinate at a much higher rate than other methods while maintaining safety.
Key Takeaway. Mixing the game’s payoff with the expert’s reference policy allows the game solver to break out of a suboptimal equilibrium.
@@ V-B Autonomous Car Racing
Reference policy and rival behaviors. To obtain a reference policy, we use a driving simulator (Fig. 6) to collect driving data (states, actions, track information, etc.) from races performed by two expert human racers in Carla [94]. With this data, we employ an inverse dynamic game training procedure inspired by works such as [38, 17]. This approach involves defining a set of basis functions for the game’s cost—each representing a distinct racing specification like optimizing lap times, collision avoidance, and track boundary adherence—and learning a context-dependent cost weight model. The model, which takes as inputs driving data and outputs the cost weights, is learned as a deep neural network leveraging automatic differentiation in JAX [92]. Given the limited volume of human-generated data, we use a dataset aggregation (DAgger) strategy [95] to enrich the training dataset for the reference policy. To obtain a stochastic reference policy, we use a game-theoretic variant of the Boltzmann distribution [9, Section 3.2], in which the Q-value function is obtained by solving the game with the inverse-learned cost weight model, which captures the noisily-rational aspect of human decision making. Finally, we produce the rival’s behavior with a defensive ILQGame policy [5], whose parameters (cost terms and weights) are not accessible to either the reference or KLGame policies.
@@
Multi-modal KLGame. Motivated by the result shown above for Basic KLGame, where blending only an aggressive reference policy can lead to unsafe outcomes, this experiment demonstrates how a multi-modal KLGame (MM-KLGame) (c.f. Sec. IV-C2) with an additional safety-enhancing mode performs in the racing scenario. The second mode we use is a following policy that controls the ego car to (i) follow the rival while keeping a constant distance when the rival is leading or (ii) defend against the rival after overtaking it. During each planning cycle, instead of picking the mode via sampling, we use rollout-based safety filters [96, Section 3.3] and perform a collision check for the KLGame-optimized trajectories associated with the overtaking mode: if they are collision-free, we choose the overtaking mode; otherwise, we pick the following mode. Again, we vary the value of $\lambda$ and inspect the qualitative difference in the resulting ego behaviors. The simulation snapshots are shown in Fig. 8. When $\lambda$ takes a relatively small value, $\lambda=5$, the MM-KLGame is not sufficiently guided by the overtaking policy and fails to overtake the rival. When $\lambda=50$, the ego using MM-KLGame behaves similarly to the unimodal case and safely overtakes the rival. However, in the case when $\lambda=5\times 10^{4}$, instead of incurring a collision as in the unimodal KLGame case, the MM-KLGame leads to a safe overtake thanks to the additional following mode in the multi-modal reference policy. We also apply the reference policy to the ego using the same mode selection rule, resulting in a shielded inverse dynamic game policy. Interestingly, the ego car fails to overtake the rival due to a more frequent application of the following policy triggered by the unsafe overtaking policy, which ultimately slows down the ego vehicle. This example also demonstrates the potential of KLGame using a multi-modal reference policy that mixes a data-driven policy and an optimization-based policy.
TABLE II: Racing Statistics
[…]
Inverse Dynamic Game [17, 38] $0.73$ $\mathbf{1.00}$
Shielded Inverse Dynamic Game $0.90$ $0.86$
KLGame (Ours) $0.85$ $0.96$
Multi-modal KLGame (Ours) $\mathbf{0.95}$ $0.92$
Racing Statistics. Finally, in order to quantitatively assess the performance of our proposed KLGame and MM-KLGame policies, we compare them against two baselines: (1). inverse dynamic game with a neural cost model (the reference policy), and (2). the shielded inverse dynamic game policy. We simulate 100 randomized racing scenarios, each with different initial positions on the Thunderhill track (for both the ego and rival vehicles), as well as varying levels of aggressiveness in the rival’s defending behaviors. Each scenario is evaluated for each policy using the same random seed. We use the same blending factor $\lambda$ for both KLGame and MM-KLGame policies across all simulation instances. The safe rate and overtaking rate statistics obtained from these trials are shown in Table II. Since the reference policy is learned from a dataset containing aggressive overtaking demonstrations, it achieves the highest overtaking rate but falls short of the safe rate. The (uni-modal) KLGame policy significantly improves the safe rate by $12\%$ compared to the reference policy, albeit at the cost of a $4\%$ reduction in the overtaking rate. The multi-modal KLGame policy, due to the additional car following mode, achieves the highest safety rate ($95\%$) among the four policies, while maintaining a decent overtaking rate ($92\%$). Finally, the shielded inverse game policy, despite reaching a high safe rate, is ultimately the most conservative policy due to the planner-agnostic safety filter.
Key Takeaway. Our blended multi-modal KLGame policy can balance between competitiveness and cautiousness in fast and rivalrous interactions such as car racing.
@@ V-C Scene-Consistent Simulated Agents in Waymax
KLGame on Waymax. For our implementation of KLGame on Waymax, we use only two generic optimization targets as cost functions: collision distance and control magnitude regularization, with the rest of the agent’s optimization objective coming from the MTR reference policy guidance. Alongside the control regularization, each agent is also tasked with maximizing its signed distance to the closest players and off-road areas. These objectives prompt the agents to prevent collisions effectively and adhere to the road. The signed distance between agents is calculated as the Minkowski difference between two footprints, while the distance to the off-road areas is obtained from road boundaries. In our experiments, all agents are replanned every second in a receding horizon fashion with a 2-second planning horizon and 0.1s discrete timestep.
Method Collision Rate $\downarrow$ ADE [3s] (m) $\downarrow$ ADE [5s] (m) $\downarrow$ ADE [8s] (m) $\downarrow$ ADE [avg] (m)
ILQGames [5] $0.13$ $0.23$ $1.02$ $3.34$ $1.53$
MaxEnt [4] $0.13$ $0.23$ $1.02$ $3.34$ $1.53$
IDM [100] $0.10$ $0.81$ $3.03$ $8.04$ $3.96$
MTR [62] $0.12$ $0.15$ $0.70$ $2.19$ $1.01$
KLGame (Ours) $0.09$ $0.18$ $0.80$ $2.45$ $1.14$
TABLE III: Results for a KLGame ego-agent planner against several other ego-agent baselines on Waymax [99] when the other agent follows the IDM [100] model.
Figure 9: As car 1 pulls out from the parked vehicle, KLGame optimizes the joint trajectories, guiding both car 0 and car 1 to exhibit yielding behaviors to prevent a collision. However, the reference MTR policy, in contrast, results in a collision.
@@
Finally, Fig. 11 shows that KLGame can yield diverse, yet scene-consistent, traffic plans, particularly when paired with a multi-modal reference policy. In this intersection scenario, agent 44 approaches the intersection from the wrong side of the road. This unlikely event leads to the marginal predictions from MTR having a hard time keeping the scenario collision-free. In the first rollout, as vehicle 44 opts for a mode to proceed straight, KLGame plans emergency maneuvers for both vehicles to avoid collision. In the subsequent rollout, vehicle 44 samples the mode of turning right. Despite a proximate interaction, KLGame computes the optimal joint trajectory, thereby successfully preventing a collision.
Waymax Statistics. We compare our proposed KLGame method against four baselines: (1) ILGames [5], (2) MaxEnt dynamic game [4], (3) the Intelligent Driver Model (IDM) [100], and (4) MTR [62]. To measure the robustness of each planner to modeling errors, we evaluated each method against an IDM agent in $500$ scenarios from the Waymo Open Motion Dataset [101] interactive validation split, which is composed of especially challenging interactive scenarios selected by [102]. The safe rate, collision rate, and ADE along each trajectory are reported in Table III. Since the ILQGames and MaxEnt dynamic game baselines do not use a multi-modal reference policy, we instead incorporate reference tracking via the most likely MTR mode. The multimodal KLGame significantly improves on the unimodal ILQGames and MaxEnt dynamic game. Furthermore, KLGame achieves the lowest collision rate among all methods, including data-driven MTR [62] while maintaining the second-lowest ADE scores by a large margin. With only a $12\%$ increase in ADE ($2.45~{}m$ with KLGame versus $2.19~{}m$ with MTR), KLGame reduces the number of collisions by $25\%$ (45 with KLGame versus 60 with MTR). Overall, these results suggest that the multi-modal policy solved by KLGame encourages safety even when the other agents in the game are not modeled perfectly.
Figure 10: As car 1 merges into the main road, the MTR-derived policy slows down car 0 but ultimately results in a collision. In comparison, KLGame optimizes the joint trajectories by planning for car 0 to switch lanes and yield, thus avoiding a collision.
Figure 11: Car 44 approaches the intersection from the incorrect side of the road. In the first rollout, car 44 opts for a mode to proceed straight, causing KLGame to plan an emergency maneuver to circumvent collision. In the subsequent rollout, car 44 samples the mode of turning right. Despite a proximate interaction, KLGame computes the optimal joint trajectory, thereby successfully averting a collision. This demonstrates KLGame’s ability to generate diverse trajectories with a multi-modal reference policy.
Key Takeaway. The KLGame planner allows integrating the state-of-the-art large-scale, data-driven behavior model with an optimization-based game policy to improve safety for complex tasks such as urban autonomous driving at scale.
Figure 12: We examined the worst case runtime and its 3-$\sigma$ limits for KLGame with different number agents. In each trial, We repeated each trails for 80 times and ensure that the KLGame solver iterates 15 times with 15 line searches.